John has been in the habit of mowing Willas lawn each week for $20. Johns opportunity cost is $15, and Willa would be willing to pay $25 to have her lawn mowed. What is the maximum tax the government can impose on lawn mowing without discouraging John and Willa from continuing their mutually beneficial arrangement?
Answer #1

John's Opportunity cost = 15$

Willa is willing to pay = 25$

John is charging = 20$

John will be happy to mowe the lawn even at 15$ as it the opportunity cost for him , which means John is ready to loose the benefit of getting 20$ for 15$.

Willa is willing to pay 25$ to John for mowing the lawn.

So the government tax rate should be in such a way that the amount that the amount John hold in hands after tax cut should not be less than 15$

So

Maximum tax amount government can impose = T = 25-15 = 10$

Maximum tax rate = (10/15)x100

Maximum tax rate = 66.67%

So government can impose a maximum tax rate of 66.67% on lawn mowing which will not discourage the relationship between John and Willa.
Answer #2

SOLUTION ;


John will not be at loss if he gets at least $15  equal to  opportunity cost. Willa is also not at a loss as he ays $25 in any case.


Hence maximum tax that can be levied

= 25 - 15

= 10 ($) (ANSWER)

answered by: Tulsiram Garg